# More is uncorrelated: Tuning the local correlations of SU($N$) Fermi-Hubbard systems via controlled symmetry breaking

### v1

@@ Abstract
Cold-atom experiments based on alkali-like atoms provide us with a tool to experimentally realize Hubbard models with a large number $N$ of components. The value of $N$ can be seen as a new handle to tune the properties of the system, leading to new physics both in the case of fully SU($N$) symmetric systems, or in the presence of controlled symmetry breaking.
We focus on the Mott transition at global half filling and we characterize local correlations between particles complementing conventional estimates with the inter-flavor mutual information. We prove that these correlations have classical nature and, using Dynamical Mean-Field Theory, we show that the SU(4) system has significantly smaller correlations than the SU(2) counterpart. In the atomic limit we prove that increasing $N$ further decreases the strength of the correlations. This suggests that a controlled reduction of the symmetry, reducing the number of effective components, can be used to enhance the degree of correlation.
We confirm this scenario solving the model for $N=4$ and gradually breaking the symmetry via a Raman field, revealing an evolution from the SU(4) to the SU(2) Mott transition as the symmetry-breaking term increases, with a sudden recovery of the large correlations of the SU(2) model at weak Raman coupling in the Mott state. By further exploring the interplay between energy repulsion and the Raman field, we obtain a rich phase diagram with three different phases – a metal, a band insulator, and a Mott insulator – all coexisting at a single tricritical point.
@@ I Introduction
Even if the fight to solve the standard spin-1/2 Hubbard model with SU(2) symmetry in two dimensions is still underway, generalizations of the model including more degrees of freedom have shown to display new fundamental physics both in a solid-state framework [2] and in quantum simulators. Some notable examples are interaction-resilient metals [3, 4, 5, 6] and selective Mott states, in which some flavors are Mott localized and others are metallic [7, 8, 9, 10]. In this context, ultracold atoms in optical lattices have flourished as powerful and flexible experimental platforms to simulate the multi-component Hubbard model, particularly leveraging the SU($N$) symmetry (associated to the nuclear spin) of alkaline-earth-like atoms such as 173Yb and 87Sr [11, 12, 13, 14, 15].
These systems also offer the possibility to drive the system from the full SU($N$) symmetry to a lower one in a continuous and controlled way. This controlled symmetry breaking has allowed to explore flavor-selective Mott localization, starting from an SU(3) model, from both the theoretical and experimental points of view [9, 10] . Flavour-selective Mott localization is indeed the extreme version of flavour-selective correlations, i.e. the simultaneous presence of particles with different degree of correlation in the same quantum system. In this perspective, it can be seen as one example of the possibility to control the degree of correlation of quantum many-body systems via the controlled breaking of the SU($N$) symmetry.
In this work we build on these ideas and we put these concepts on more solid grounds assessing the strength and the nature of the correlation properties as the number of components $N$ varies.
@@
This program exploits the development of a unified geometrical framework [55] for the quantification and classification of resources in open quantum systems, that allows to evaluate the informational distance of a given density matrix from reference gaussian states [48, 49, 47], separable states [35, 34] or pseudo-classical states [39, 40, 38].
@@
Here we leverage on this result to show that for multi-component fermionic models that conserve the number of particles in each flavor, such as the SU($N$) Hubbard model, the local correlations are effectively classical and they are not related to local entanglement, nor quantum discord. Hence, these correlations are associated, at the local level, with pure nongaussian resources.
By studying the half-filled SU($N$) Hubbard model (i.e. with $N/2$ fermions per site) by means of Dynamical Mean-Field Theory (DMFT) [57], we show that these classical correlations are dramatically reduced going from the popular SU(2) model to SU(4). Solving exactly the atomic limit, we provide an evidence that this trend is more general, with the strength of the correlations decreasing with the number of components $N$, leading to an effective decoupling in the large-$N$ limit.
We finally show that a controlled breaking of the SU(4) symmetry provides a tool to effectively reduce the number of components, hence increasing the correlations. This scenario can be experimentally realized and detected in cold-atoms based simulators of SU($N$) fermions.
After introducing the model in Sec. II, and the entropy-related quantities that we use to quantify correlations in Sec. III, we consider the SU(4)-symmetric Hubbard model, comparing it to the SU(2) case, in Sec. IV.1. By means of inter-flavor mutual information, we show that correlations in the Mott phase are significantly weaker for the four-component system than for the two-component one. We then extend this result analytically in Sec. IV.2, by studying the atomic limit of an SU($N$)-symmetric Hubbard model, demonstrating that mutual information vanishes in the large-$N$ limit, leading to an uncorrelated Mott insulator. Finally, in Sec. IV.3 we explore the effects of symmetry breaking, introducing a term that lifts the degeneracy between two of the four flavors. This induces a crossover from SU(4) to SU(2), tunable via a symmetry-breaking term.
@@ III Assessing correlations through entropic quantities
In this work, we specify these ideas to our multi-component system by focusing on the local inter-flavor mutual information, a rather general measure of correlations between two out of the $N$ flavors of the model. For the SU(2) case, this quantity has already been introduced for the two spin species as intra-orbital mutual information and it has been found to faithfully account for the correlation properties of Hubbard systems [56, 30]. Additionally, for systems where each fermionic flavor is individually conserved, the inter-flavor mutual information depends only on experimentally accessible quantities, such as average double occupancies and flavor-resolved densities.
@@ III.1 Two-flavor and single-flavor entropy
The main target of our analysis are the local correlations between pairs of flavors. Therefore, we start from the local two-flavor reduced density matrix $\rho^{(i)}_{\alpha\beta}\equiv\rho_{\alpha\beta}$, obtained tracing out all lattice sites except site $i$ and all flavors different from $\alpha$ and $\beta$. In the absence of inhomogeneities and spatial symmetry breaking, we can drop the lattice site label $i$. From this we compute the two-flavor entropy
@@
where $p_{0},p_{\alpha},p_{\beta},p_{\alpha,\beta}$ represent the probabilities of the site being empty, singly occupied by flavor $\alpha$, singly occupied by flavor $\beta$, and occupied by both flavors, respectively [28, 79, 30], regardless of the other flavours. One has $p_{\alpha,\beta}=\langle n_{\alpha}n_{\beta}\rangle$, $p_{\alpha}=\langle n_{\alpha}\rangle-\langle n_{\alpha}n_{\beta}\rangle$, $p_{\beta}=\langle n_{\beta}\rangle-\langle n_{\alpha}n_{\beta}\rangle$ and $p_{0}=1-\langle n_{\alpha}\rangle-\langle n_{\beta}\rangle+\langle n_{\alpha}n_{\beta}\rangle$. Hence, the two-flavor density matrix, as well as all the entropic quantities that depend on it, can be easily computed in terms of observables accessible in cold atom experiments [10], providing a much simpler measure of correlations with respect to those based on the knowledge of the full $N$-flavor density matrix.
@@ IV Results
In this section, we present our DMFT results. In Sec. IV.1, we focus on the SU(4) symmetric Hubbard model, highlighting its differences and analogies with the SU(2) model. In order to further explore the role of the number of components, in Sec. IV.2 we study the SU($N$) symmetric model in the atomic limit, where we can vary $N$ easily. Finally, in Sec. IV.3, we study the breaking the SU(4) symmetry of Eq. 1 by allowing $\Omega\neq 0$.
@@ IV.1 SU(4) vs SU(2) Mott transitions
Figure 3: Two-flavor entropy $s_{12}$, single-flavor entropy $s_{1}=s_{2}$ and inter-flavor mutual information $I_{12}$, plotted against the interaction strength, for the SU($2$) Hubbard model (top panel) and the SU($4$) model (bottom panel). The area under the mutual information is shaded to indicate that all local correlators are contained beneath it.  All quantities are in units of $\log(2)$ (bits, in information theory language).
@@
with $\Sigma_{\alpha\alpha}(i\omega)$ the momentum-independent diagonal self-energy matrix. Here, because of the SU($N$) symmetry, $Z_{\alpha}$ does not depend on $\alpha$. $Z_{\alpha}$ measures the weight of the quasiparticle excitations, and it equals 1 in a noninteracting metal, while it vanishes in a Mott-Hubbard insulating state. Our results, shown in Fig. 2, confirm previous literature and find a continuous evolution of $Z_{\alpha}$ before reaching a critical interaction $U/D\simeq 2.8$ for the SU(2) case and a significantly larger value of $U/D\simeq 4.2$ for SU(4). Besides this quantitative difference, the two transitions seem similar based on these results.
@@
The mutual information between two flavors, presented in Fig. 3, reveals instead a completely different picture in SU(2) and SU(4) systems. Indeed, while the SU(2) case has $I_{12}$ continuously approaching a value close to $\log{2}$ at the Mott transition, the same quantity appears always small for the SU(4) system. Yet, if we zoom in, we see that also in this case $I_{12}$ is a monotonically increasing function of $U/t$, but it reaches a maximum value of $0.08\log{2}$, signaling significantly lower correlations in the Mott insulator compared to the SU(2) case. The much smaller value descends from the fact that $s_{12}$ remains close to $2\log{2}$ as we increase $U/t$, in contrast with the SU(2) case, where it drops from $2\log{2}$ to $\log{2}$. Our results show that the behavior of $Z$ hides a profound difference between the SU(2) and the SU(4) systems. In the next section, we discuss the effect of further increasing $N$ in the atomic limit, where the model can be solved for arbitrary $N$.
IV.2 Atomic SU($N$) model at any value of NN
@@
Figure 4: Inter-flavor mutual information (top) and double occupancy per site (bottom) for a half-filled SU($N$) Hubbard model in the atomic limit $t=0$, plotted against the total number of flavors NN. In the inset we show again the mutual information vs. NN in a log-log plot highlighting the $1/2N^{2}$ scaling (red dashed line).
@@
Figure 5: Double occupancies per site for the Raman-free flavors (top panel) and the Raman-coupled flavors (bottom panel) for different values of the Raman coupling $\Omega$. For Raman-coupled flavors, double occupancies vanish for $U>U^{\mathrm{R}}_{\mathrm{c}}(\Omega)$ due to the occurrence of complete polarization, while the Raman-free ones display a crossover between the values in the SU($4$) and SU($2$) Hubbard model.
@@
Figure 6: Quasiparticle weight for the Raman-free flavors (top panel) and the Raman-coupled flavors (bottom panel) for different values of the Raman coupling $\Omega$. For Raman-coupled flavors, the quasiparticle weight tends to $1$, signaling the hybrid nature of the insulating state, while the Raman-free $Z$ vanishes for $U>U^{\mathrm{M}}_{\mathrm{c}}(\Omega)$, marking the onset of the Mott metal-to-insulator transition.
@@
In order to extend results to large values of $N$, for which a DMFT calculation would rapidly become unfeasible, we consider the atomic limit $t=0$ where the model simplifies to a collection of disconnected Hubbard atoms. The density of each flavor is $\langle n_{\alpha}\rangle=n/N$, where $n$ is the number of fermions per site, while the double occupancies are given by the ratio between all possible couples among $n$ particles and all possible couples among $N$ flavors, i.e. $\langle n_{\alpha}n_{\beta}\rangle=\frac{n(n-1)}{N(N-1)}$ for every $\alpha\neq\beta$.
@@
In Fig. 4 we display how the inter-flavor mutual information and the double occupancies evolve with the number of flavors at half filling ($n=N/2$). The mutual information decreases rapidly with the number of flavors, leading to very weakly correlated Mott insulators as $N$ becomes large. Hence, in this limit the Mott state, despite being driven by large interactions, lacks inter-flavor correlations. Consequently, since the mutual information is an upper bound to all correlation functions, Eq. 9 implies that, in the large-$N$ limit, the correlation function of any pair of local operators must vanish at least as fast as $I_{\alpha\beta}$. , approaching an uncorrelated mean-field description, with $\langle\mathcal{O}_{\alpha}\mathcal{O}_{\beta}\rangle\to\langle\mathcal{O}_{\alpha}\rangle\langle\mathcal{O}_{\beta}\rangle,\,\forall\mathcal{O}_{\alpha},\mathcal{O}_{\beta}$. We show the case of double occupancies, that approach their factorized value $\langle n_{\alpha}n_{\beta}\rangle\to\langle n_{\alpha}\rangle\langle n_{\beta}\rangle=1/4$ for $N\rightarrow\infty$, as shown in the bottom panel of Fig. 4. Our analytical treatment gives thus a strong indication that, in the atomic limit, the local correlations approach a mean field behavior, where the flavors are decoupled. Furthermore, at half filling it can be analytically proven that the leading term for the mutual information at large $N$ scales as $1/2N^{2}$, as confirmed by the inset of Fig. 4. It is interesting to observe that the drop from the large value of $N=2$ to small values is very rapid, and already $N=4$ has a remarkably smaller value of $I_{12}$. In this light the standard two-species Mott transition appears to be intrinsically different from any other value of $N$. We finally mention that the trends are similar also for any other integer filling $n\neq N/2$.

### v2

# More is uncorrelated: Tuning the local correlations of SU(NN) Fermi-Hubbard systems via controlled symmetry breaking

@@ Abstract
Cold-atom experiments based on alkali-like atoms provide us with a tool to experimentally realize Hubbard models with a large number NN of components. The value of NN can be seen as a new handle to tune the properties of the system, leading to new physics both in the case of fully SU(NN) symmetric systems, or in the presence of controlled symmetry breaking. We focus on the Mott transition at global half filling and we characterize local correlations between particles through the inter-flavor mutual information, an experimentally accessible quantity that rigorously measures the distance from the closest gaussian state, unveiling features that cannot be accessed by conventional probes of Mottness. We prove that these correlations are fully independent from local entanglement and quantum discord, and, using Dynamical Mean-Field Theory, we show that the SU(4) system has significantly smaller correlations than the SU(2) counterpart. In the atomic limit we prove that increasing NN further decreases the strength of the correlations. This suggests that a controlled reduction of the symmetry, reducing the number of effective components, can be used to enhance the degree of correlation. We confirm this scenario solving the model for $N=4$ and gradually breaking the symmetry via a Raman field, revealing an evolution from the SU(4) to the SU(2) Mott transition as the symmetry-breaking term increases, with a sudden recovery of the large correlations of the SU(2) model at weak Raman coupling in the Mott state. By further exploring the interplay between energy repulsion and the Raman field, we obtain a rich phase diagram with three different phases – a metal, a band insulator, and a Mott insulator – all coexisting at a single tricritical point.
@@ I Introduction
Even if the fight to solve the standard spin-1/2 Hubbard model with SU(2) symmetry in two dimensions is still underway, generalizations of the model including more degrees of freedom have shown to display new fundamental physics both in a solid-state framework [2] and in quantum simulators. Some notable examples are interaction-resilient metals [3, 4, 5, 6] and selective Mott states, in which some flavors are Mott localized and others are metallic [7, 8, 9, 10]. In this context, ultracold atoms in optical lattices have flourished as powerful and flexible experimental platforms to simulate the multi-component Hubbard model, particularly leveraging the SU(NN) symmetry (associated to the nuclear spin) of alkaline-earth-like atoms such as 173Yb and 87Sr [11, 12, 13, 14, 15].
These systems also offer the possibility to drive the system from the full SU(NN) symmetry to a lower one in a continuous and controlled way. This controlled symmetry breaking has allowed to explore flavor-selective Mott localization, starting from an SU(3) model, from both the theoretical and experimental points of view [9, 10]. Flavor-selective Mott localization is indeed the extreme version of flavor-selective correlations, i.e. the simultaneous presence of particles with different degree of correlation in the same quantum system. In this perspective, it can be seen as one example of the possibility to control the degree of correlation of quantum many-body systems via the controlled breaking of the SU(NN) symmetry.
In this work we build on these ideas and we put these concepts on more solid grounds assessing the strength and the nature of the correlation properties as the number of components NN varies.
@@
This program exploits the development of a unified geometrical approach [55] for the quantification and classification of resources in open quantum systems, that allows to evaluate the informational distance of a given density matrix from reference gaussian states [48, 49, 47], separable states [35, 34] or pseudo-classical states [39, 40, 38].
@@
Here we leverage on this result to show that for multi-component fermionic models that conserve the number of particles in each flavor, such as the SU(NN) Hubbard model, the local correlations are effectively classical and they are not related to local entanglement, nor quantum discord. Hence, these correlations are associated, at the local level, with pure nongaussian resources.
By studying the half-filled SU(NN) Hubbard model (i.e. with $N/2$ fermions per site) by means of Dynamical Mean-Field Theory (DMFT) [57], we show that these classical correlations are dramatically reduced going from the popular SU(2) model to SU(4). Solving exactly the atomic limit, we provide an evidence that this trend is more general, with the strength of the correlations decreasing with the number of components NN, leading to an effective decoupling in the large-NN limit.
We finally show that a controlled breaking of the SU(4) symmetry provides a tool to effectively reduce the number of components, hence increasing the correlations. This scenario can be experimentally realized and detected in cold-atoms based simulators of SU(NN) fermions.
After introducing the model in Sec. II, and the entropy-related quantities that we use to quantify correlations in Sec. III, we consider the SU(4)-symmetric Hubbard model, comparing it to the SU(2) case, in Sec. IV.1. By means of inter-flavor mutual information, we show that correlations in the Mott phase are significantly weaker for the four-component system than for the two-component one. We then extend this result analytically in Sec. IV.2, by studying the atomic limit of an SU(NN)-symmetric Hubbard model, demonstrating that mutual information vanishes in the large-NN limit, leading to an uncorrelated Mott insulator. Finally, in Sec. IV.3 we explore the effects of symmetry breaking, introducing a term that lifts the degeneracy between two of the four flavors. This induces a crossover from SU(4) to SU(2), tunable via a symmetry-breaking term.
@@ III Assessing correlations through entropic quantities
In this work, we specify these ideas to our multi-component system by focusing on the local inter-flavor mutual information, a rather general measure of correlations between two out of the NN flavors of the model. For the SU(2) case, this quantity has already been introduced for the two spin species as intra-orbital mutual information and it has been found to faithfully account for the correlation properties of Hubbard systems [56, 30]. Additionally, for systems where each fermionic flavor is individually conserved, the inter-flavor mutual information depends only on experimentally accessible quantities, such as average double occupancies and flavor-resolved densities.
@@ III.1 Two-flavor and single-flavor entropy
The main target of our analysis are the local correlations between pairs of flavors. Therefore, we start from the local two-flavor reduced density matrix $\rho^{(i)}_{\alpha\beta}\equiv\rho_{\alpha\beta}$, obtained tracing out all lattice sites except site $i$ and all flavors different from $\alpha$ and $\beta$. In the absence of inhomogeneities and spatial symmetry breaking, we can drop the lattice site label ii. From this we compute the two-flavor entropy
@@
where $p_{0},p_{\alpha},p_{\beta},p_{\alpha,\beta}$ represent the probabilities of the site being empty, singly occupied by flavor $\alpha$, singly occupied by flavor $\beta$, and occupied by both flavors, respectively [28, 79, 30], regardless of the other flavors. One has $p_{\alpha,\beta}=\langle n_{\alpha}n_{\beta}\rangle$, $p_{\alpha}=\langle n_{\alpha}\rangle-\langle n_{\alpha}n_{\beta}\rangle$, $p_{\beta}=\langle n_{\beta}\rangle-\langle n_{\alpha}n_{\beta}\rangle$ and $p_{0}=1-\langle n_{\alpha}\rangle-\langle n_{\beta}\rangle+\langle n_{\alpha}n_{\beta}\rangle$. Hence, the two-flavor density matrix, as well as all the entropic quantities that depend on it, can be easily computed in terms of observables accessible in cold atom experiments [10], providing a much simpler measure of correlations with respect to those based on the knowledge of the full NN-flavor density matrix.
@@ IV Results
In this section, we present our DMFT results. In Sec. IV.1, we focus on the SU(4) symmetric Hubbard model, highlighting its differences and analogies with the SU(2) model. In order to further explore the role of the number of components, in Sec. IV.2 we study the SU(NN) symmetric model in the atomic limit, where we can vary NN easily. Finally, in Sec. IV.3, we study the breaking the SU(4) symmetry of Eq. 1 by allowing $\Omega\neq 0$.
@@ IV.1 SU(4) vs SU(2) Mott transitions
Figure 3: Two-flavor entropy $s_{12}$, single-flavor entropy $s_{1}=s_{2}$ and inter-flavor mutual information $I_{12}$, plotted against the interaction strength, for the SU(22) Hubbard model (top panel) and the SU(44) model (bottom panel). The area under the mutual information is shaded to indicate that all local correlators are contained beneath it.  All quantities are in units of $\log(2)$ (bits, in information theory language).
@@
with $\Sigma_{\alpha\alpha}(i\omega)$ the momentum-independent diagonal self-energy matrix. Here, because of the SU(NN) symmetry, $Z_{\alpha}$ does not depend on $\alpha$. $Z_{\alpha}$ measures the weight of the quasiparticle excitations, and it equals 1 in a noninteracting metal, while it vanishes in a Mott-Hubbard insulating state. Our results, shown in Fig. 2, confirm previous literature and find a continuous evolution of $Z_{\alpha}$ before reaching a critical interaction $U/D\simeq 2.8$ for the SU(2) case and a significantly larger value of $U/D\simeq 4.2$ for SU(4). Besides this quantitative difference, the two transitions seem similar based on these results.
@@
The mutual information between two flavors, presented in Fig. 3, reveals instead a completely different picture in SU(2) and SU(4) systems. Indeed, while the SU(2) case has $I_{12}$ continuously approaching a value close to $\log{2}$ at the Mott transition, the same quantity appears always small for the SU(4) system. Yet, if we zoom in, we see that also in this case $I_{12}$ is a monotonically increasing function of $U/t$, but it reaches a maximum value of $0.08\log{2}$, signaling significantly lower correlations in the Mott insulator compared to the SU(2) case. The much smaller value descends from the fact that $s_{12}$ remains close to $2\log{2}$ as we increase $U/t$, in contrast with the SU(2) case, where it drops from $2\log{2}$ to $\log{2}$. Our results show that the behavior of $Z$ hides a profound difference between the SU(2) and the SU(4) systems. In the next section, we discuss the effect of further increasing NN in the atomic limit, where the model can be solved for arbitrary NN.
IV.2 Atomic SU(NN) model at any value of NN
@@
Figure 4: Inter-flavor mutual information (top) and double occupancy per site (bottom) for a half-filled SU(NN) Hubbard model in the atomic limit $t=0$, plotted against the total number of flavors NN. In the inset we show again the mutual information vs. NN in a log-log plot highlighting the $1/2N^{2}$ scaling (red dashed line).
@@
Figure 5: Double occupancies per site for the Raman-free flavors (top panel) and the Raman-coupled flavors (bottom panel) for different values of the Raman coupling $\Omega$. For Raman-coupled flavors, double occupancies vanish for $U>U^{\mathrm{R}}_{\mathrm{c}}(\Omega)$ due to the occurrence of complete polarization, while the Raman-free ones display a crossover between the values in the SU(44) and SU(22) Hubbard model.
@@
Figure 6: Quasiparticle weight for the Raman-free flavors (top panel) and the Raman-coupled flavors (bottom panel) for different values of the Raman coupling $\Omega$. For Raman-coupled flavors, the quasiparticle weight tends to 11, signaling the hybrid nature of the insulating state, while the Raman-free $Z$ vanishes for $U>U^{\mathrm{M}}_{\mathrm{c}}(\Omega)$, marking the onset of the Mott metal-to-insulator transition.
@@
In order to extend results to large values of NN, for which a DMFT calculation would rapidly become unfeasible, we consider the atomic limit $t=0$ where the model simplifies to a collection of disconnected Hubbard atoms. The density of each flavor is $\langle n_{\alpha}\rangle=n/N$, where $n$ is the number of fermions per site, while the double occupancies are given by the ratio between all possible couples among $n$ particles and all possible couples among NN flavors, i.e. $\langle n_{\alpha}n_{\beta}\rangle=\frac{n(n-1)}{N(N-1)}$ for every $\alpha\neq\beta$.
@@
In Fig. 4 we display how the inter-flavor mutual information and the double occupancies evolve with the number of flavors at half filling ($n=N/2$). The mutual information decreases rapidly with the number of flavors, leading to very weakly correlated Mott insulators as NN becomes large. Hence, in this limit the Mott state, despite being driven by large interactions, lacks inter-flavor correlations. Consequently, since the mutual information is an upper bound to all correlation functions, Eq. 9 implies that, in the large-NN limit, the correlation function of any pair of local operators must vanish at least as fast as $I_{\alpha\beta}$. , approaching an uncorrelated mean-field description, with $\langle\mathcal{O}_{\alpha}\mathcal{O}_{\beta}\rangle\to\langle\mathcal{O}_{\alpha}\rangle\langle\mathcal{O}_{\beta}\rangle,\,\forall\mathcal{O}_{\alpha},\mathcal{O}_{\beta}$. We show the case of double occupancies, that approach their factorized value $\langle n_{\alpha}n_{\beta}\rangle\to\langle n_{\alpha}\rangle\langle n_{\beta}\rangle=1/4$ for $N\rightarrow\infty$, as shown in the bottom panel of Fig. 4. Our analytical treatment gives thus a strong indication that, in the atomic limit, the local correlations approach a mean field behavior, where the flavors are decoupled. Furthermore, at half filling it can be analytically proven that the leading term for the mutual information at large NN scales as $1/2N^{2}$, as confirmed by the inset of Fig. 4. It is interesting to observe that the drop from the large value of $N=2$ to small values is very rapid, and already $N=4$ has a remarkably smaller value of $I_{12}$. In this light the standard two-species Mott transition appears to be intrinsically different from any other value of NN. We finally mention that the trends are similar also for any other integer filling $n\neq N/2$.